OpenStudy (connor):

A man built a house costing 9 times as much as the lot on which it was built. If the lot cost $5000 less than 1/6 the cost of the house. What was the cost of each?

OpenStudy (anonymous):

Price of lot is 1/9 the price of house. \[1/9 = 1/6 -5000\] \[5000 + 1/9 = 1/6\] \[1/6-1/9=5000\] \[1/18 = 5000\] Therefore 2/18 = 1/9 = 10000 = Price of lot Price of house = 10000*9= 90000

OpenStudy (anonymous):

Solve the following equations for house and lot:\[\left\{\text{house} \text{ = } 9*\text{lot},\text{ lot }\text{=} \frac{1}{6}\text{ house}-5000\right\}\]\[\{\text{lot}=\text{$\$$10000},\text{ house}=\text{$\$$90000}\} \]